# Sub-keV dark matter can strongly ionize molecular clouds

### v1

@@ IV Results
These model-independent constraints for DM producing two photons can be directly expressed as constraints in the coupling of different feebly interacting particles, such as ALPs, scalars or sterile neutrinos. We calculate our constraint on the coupling of ALPs with photons as Carenza et al. (2025)
where is the model independent limit on the DM lifetime and ma is the ALP mass, which is from 10 eV to 10 keV in our analysis. The emergence of ALPs in the eV-keV mass range might be unrelated to solving the strong CP problem, and for this reason we refer to them as ALPs and not QCD axions Peccei (2008); Choi et al. (2021); Gendler et al. (2024); Peccei and Quinn (1977); Weinberg (1978); Wilczek (1978). Several QCD axion and ALP models predict viable DM in the eV–keV range, motivating searches in this regime Di Luzio et al. (2021); Giannotti (2005); Di Luzio et al. (2020); Rubakov (1997); Gavela et al. (2024); de Giorgi and Ramos (2025); O’Hare et al. (2022); Gelmini et al. (2021); Ayala et al. (2014); Carenza et al. (2020); Lucente et al. (2022).
In Fig. 4 we report our bounds for the coupling of ALPs with photons, gaγ, for keV mass ALPs, compared to previous constraints. We highlight here that, even in the most pessimistic case of L1551, these constraints are competitive with constraints from reionization Cadamuro and Redondo (2012) and extend to higher masses. For our fiducial constraint, from the DRAGON cloud, we significantly improve on the CMB and LeoT constraints up to several hundreds of eV. We also point out that constraints based on observations of Leo T are influenced by different sources of uncertainty in their model of the gas distribution, their temperature and its DM halo Wadekar and Wang (2022), that are expected to be similar to those affecting the ionization rate measurement for the DRAGON cloud. The halo model used to derive their constraint is discussed in Ref. Faerman et al. (2013), whose uncertainties can mildly affect the resulting constraint. However, a more important issue is connected to the possible breaking of the hydrostatic equilibrium assumption, fundamental for the model of Ref. Faerman et al. (2013). Indeed, even though observations do not point to a rotation of the HI gas in Leo T, it is possible that the gas is rotating in a direction aligned with our view, which would conceal its motion. However, Ref. Wadekar and Wang (2022) claimed that these uncertainties are under control and the failure of hydrostatic equilibrium might even lead to a stronger bound. Moreover, our optimistic forecast for G1.4-1.8+87, that represents what we expect from a low ionization cloud near the Galactic center, illustrates that MC observations could be the leading constraints in the sub-keV range for ALPs and similar particles.
@@
We also report the constraint for scalar particles as DM, that have mixing angle with the SM Higgs Dev et al. (2021); Nguyen et al. (2024). We calculate this mixing angle from the scalar dark matter decay to two photons lifetime as
@@
where is the fine structure constant, GeV is the electroweak vacuum expectation value and mS is the mass of the scalar particle. A comparison of our constraints with previous ones, for scalar DM, is shown in Fig. 5. As for the case of ALPs, we are able to improve on existent constraints, such as those from CMB spectral distortions Bolliet et al. (2021), LeoT gas heating Wadekar and Farrar (2021) or the model independent constraint on decaying DM (dDM) derived in Refs. Audren et al. (2014); Poulin et al. (2016), and our results are complementary with X-ray constraints Gewering-Peine et al. (2017); Essig et al. (2013); Foster et al. (2021), constraints from the CMB anisotropies Slatyer (2016) and stellar constraints Dev et al. (2021); Balaji et al. (2022). Interestingly, MCs can probe a region of the parameter space that cannot be currently tested from arguments of the stars dynamics.
@@
where m is the sterile neutrino mass and GF is the Fermi constant that represents the strength of the weak force. In this case, instead of producing two photons in the decay, only one is produced. Therefore, the value adopted is a factor of 2 lower than the one shown in Fig. 3. These bounds are shown in Fig. 6, compared to previous ones. As one can see, the parameter space that is covered by MCs is already well tested by other mechanisms, specially the Lyman-alpha forest Bringmann et al. (2023). Also the Tremaine-Gunn constraint Tremaine and Gunn (1979) rules out sterile neutrinos with masses below keV. We report the optimistic constraint derived from G1.4-1.8+87, which may become relevant again if model-dependent approaches reopen this region of parameter space Dev et al. (2025).
@@
Finally, we have also calculated these constraints for the case of vector bosons decaying into three photons. In this case, the spectrum of photons is given by a continuum spectrum, instead of a line signal. The photon flux at a given position can be expressed as Pospelov et al. (2008); Linden et al. (2024a); McDermott et al. (2018); Cheung et al. (2025); Nguyen (2025); Redondo and Postma (2009); Krnjaic et al. (2023); Linden et al. (2024b)
@@
where . Inserting this term in Eq. 4, we can compute Jχ→3γ as in Eq. 3, for three-photon final state. The result is shown in Fig. 7, upper panel. A trend similar to the monochromatic injection appears, with a larger ionization at lower DM masses. In Fig. 7 (lower panel) we show the limit on the inverse decay rate of three-photon final state injection, for the case of L1551 and G1.4-1.8+87.
@@
We derive constraints on well-motivated vector dark matter models, such as the dark photon and the vector boson, using decay rate limits of the vector-to-three-photon final state. Both models arise from extensions of the Standard Model by anomaly-free gauge symmetries. The dark photon interacts with Standard Model particles via kinetic mixing with the photon field Holdom (1986); Fayet (1980, 1981); Kahn et al. (2017); Fabbrichesi et al. (2020); Caputo et al. (2021), while the vector boson couples directly to the Standard Model through a flavor-universal coupling Chun and Yun (2022); Carena et al. (2004); Nguyen and Tait (2023); Smith and Tait (2024). These symmetry extensions are extensively studied in the contexts of string theory, string compactification, and vector dark matter production via quantum fluctuations during inflation Servant and Tait (2003); Graham et al. (2016); Hebecker et al. (2024); Cyncynates and Weiner (2025a, b).
@@
Figure 8 presents the optimistic G1.4-1.8+87 constraints on vector dark matter models. We compute the decay width using the Euler-Heisenberg (EH) approximation, which is valid in the keV mass range McDermott et al. (2018); Linden et al. (2024b). The top panel shows the constraint on kinetic mixing , compared to limits from the universe’s age (4.6 Gyr), x-ray observations by INTEGRAL Linden et al. (2024b), and direct detection bounds from XENON Aprile et al. (2019). For vector dark matter, we derive constraints on the flavor-universal coupling , assuming the decay proceeds dominantly via electron loops under the EH approximation. We compare this result with the INTEGRAL limit on the same three-photon final state Linden et al. (2024b) and with theoretical constraints based on neutrino decay channels and the universe’s lifetime Chun and Yun (2022).
@@
V Discussion and conclusions
In this paper, we investigate the ionization of neutral gas produced in dense MCs by decaying/annihilating DM. In particular, sub-keV DM could generate UV/X-ray photons that are able to produce a high and homogeneous ionization across the cloud since they are absorbed efficiently within such systems. In fact, we show that the ionization produced by the photon emission from sub-keV DM saturates the ionization that is observed in different MCs for decay lifetimes, or annihilation rates, that are not ruled out by any other constraint. Therefore, we test for the first time MCs as a target for indirect DM searches, demonstrating that they allow us to improve current constraints for decaying DM particles such as ALPs, dark photons, scalar particles or sterile neutrinos for masses from a few tens of eV to slightly below keV.
Interestingly, The ionization of dense MCs is expected to be primarily driven by CRs, as UV light cannot effectively penetrate these regions. However, sub-keV unstable DM particles could be producing UV radiation directly inside the MC, leading to a persistent background of ionization, potentially influencing both the star formation rates and the chemical evolution within the clouds. Notably, a recent study suggests that DM particles with MeV-scale masses might account for the anomalously high ionization rates observed in the Galactic Center De la Torre Luque et al. (2025).
We begin by identifying local MCs as a particularly robust probe of this kind of DM. This is because estimating their ionization rate requires only knowledge of the local DM density — without the need to assume a specific DM distribution model. Using the local MC L1551, which is not the cloud with the lowest observed ionization rate, we derive constraints that are similar, or even better, than the previous leading constraints. Many of the existing bounds rely on assumptions that can affect their reliability and expose them to significant uncertainties, as opposed to the more robust MC bounds that we explored. Then, we examine the case of G28.37+00.07, or the DRAGON cloud, which has been studied previously in detail and shows regions of it that are exceptionally quiescent in terms of their star formation activity, with ionization rates compatible with s-1. This constitutes what we call our fiducial constraint, and allows us improve previous constraints from CMB and LeoT gas heating up to several hundreds of eV. Finally, we use the case of G1.4-1.8+87, an MC that is pc away from the Galactic Center, to show the potential of using clouds near the Galactic Center. For this case, we take the values obtained in Ref. Bhoonah et al. (2019) to illustrate that, assuming a NFW DM density profile, we reach the strongest DM bounds to date in a mass range from a few tens of eV to the keV. However, this must be taken as an optimistic forecast since the characterization of the temperature of that could was put in doubt in Ref. Farrar et al. (2020). We highlight that clouds that are closer to the Galactic Center can be found, therefore providing even stronger DM bounds. Finally, we highlight that the most promising targets for indirectly searching for sub-keV DM decays or annihilations into photons are IRDCs located near the Galactic Center but slightly above the Galactic plane. These regions are expected to be less affected by CR ionization, which is sourced in the Galactic disk. As a result, such clouds are likely to exhibit very low ionization levels, making any DM-induced effects more distinguishable. Although a more comprehensive survey of MCs will further refine the constraints presented here, our results already demonstrate the strong potential of this approach for probing different types of light DM particles. We also note that this strategy can be also applied for MCs in nearby galaxies.
Among the expected future refinements, we list a few below.
i) A more exhaustive survey of the measurements of dark MCs, dedicated to identify clouds with lower ionization fractions. These measurements are often complex, and they are usually focused on measuring the ionization around clouds with regions of star formation. However, we show that dark clouds constitute a very interesting target for DM searches, hoping to increase the interest towards targeting dark MCs for observations. In addition, if dark clouds systems can be found closer to the GC, our limits will be highly improved. In this regard, there are already extensive surveys of such clouds, like the analyses from Ref. Simon et al. (2006). A dedicated exploration and analysis could yield improvements of up to a few orders of magnitude in the derived DM constraints.
iii) Probably the most obvious improvement would be that of modeling the ionization produced by CRs. Provided that CRs are expected to be the main ionization agent of dense MCs, including their contribution must certainly improve our constraints. Unfortunately, this estimation is not simple and requires an accurate knowledge of the distribution of CRs in different regions of the Galaxy and at very low energies, where our models are not robust enough to produce a reliable constraint. However, reasonable estimations can be made for local MCs, and improvements of a factor of a few in the constraints can be already obtained with conservative assumptions, as discussed above.
As a final point, we also note that X-ray telescopes are very promising for constraining DM above a few hundreds eV. Especially eROSITA could place constraints above keV that can beat all previous constraints Dekker et al. (2021), once they have enough observational time. However, the the current constraints from eROSITA or Chandra below keV are even weaker than those that we obtain and forecast in this work.
In conclusion, we have shown the potential impact of sub-keV DM decaying/annihilating into photons in dense gas regions, which can be specially relevant for probing ALPs or scalar DM particles. Our results demonstrate that astrochemical observations in MCs can serve as a novel and powerful probe for indirect searches of sub-keV DM. This approach offers particularly robust constraints across a range of particle physics models and has the potential to surpass existing bounds for DM masses between approximately eV and keV.
VI Acknowledgements
We thank Torsten Bringmann for fruitful discussion. PDL is supported by the Juan de la Cierva JDC2022-048916-I grant, funded by MCIU/AEI/10.13039/501100011033 European Union ”NextGenerationEU”/PRTR. The work of PDL is also supported by the grants PID2021-125331NB-I00 and CEX2020-001007-S, both funded by MCIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”. PDL also acknowledges the MultiDark Network, ref. RED2022-134411-T. This project used computing resources from the National Academic Infrastructure for Supercomputing in Sweden (NAISS) under project NAISS 2024/5-666. PC and TTQN are supported by the Swedish Research Council under contract 2022-04283. This article/publication is based on the work from COST Action COSMIC WISPers CA21106, supported by COST (European Cooperation in Science and Technology).

### v2

@@ Abstract
These model-independent constraints for DM producing two photons can be directly expressed as constraints in the coupling of different feebly interacting particles, such as ALPs, scalars or sterile neutrinos. We calculate our constraint on the coupling of ALPs with photons as Carenza et al. (2025b)
where is the model independent limit on the DM lifetime and ma is the ALP mass.
In Fig. 1 we report bounds for the ALPs-photon coupling, gaγ, for keV mass ALPs, compared to previous constraints. Even in the most pessimistic case of L1551, these constraints are competitive with constraints from reionization Cadamuro and Redondo (2012) and extend to higher masses. For our fiducial constraint, from the DRAGON cloud, we significantly improve on the CMB and LeoT constraints up to several hundreds of eV. Moreover, our optimistic forecast for G1.4-1.8+87, that represents what we expect from a low ionization cloud near the GC, illustrates that MC observations could be the a significant constraints in the sub-keV range for ALPs and similar particles. We also provide constraints for other well-motivated DM models in the Supplementary Material. This Supp. Material contains the following references: Yan et al. (2001); Burkert (1995); Moore and others (1999); Cirelli et al. (2011); Di Mauro (2021); Dev et al. (2021); Nguyen et al. (2024); Bolliet et al. (2021a); Wadekar and Farrar (2021); Audren et al. (2014); Poulin et al. (2016); Gewering-Peine et al. (2017); Essig et al. (2013); Foster et al. (2021); Slatyer (2016); Fiorillo et al. (2025); Bottaro et al. (2023); Bolliet et al. (2021b); Boyarsky et al. (2019); Wadekar and Wang (2022); Bezrukov et al. (2010); Bringmann et al. (2023); Tremaine and Gunn (1979); Dev et al. (2025); Pospelov et al. (2008); Linden et al. (2024a); McDermott et al. (2018); Cheung et al. (2025); Nguyen (2025); Redondo and Postma (2009); Krnjaic et al. (2023); Linden et al. (2024b); Holdom (1986); Fayet (1980, 1981); Kahn et al. (2017); Fabbrichesi et al. (2020); Caputo et al. (2021); Chun and Yun (2022); Carena et al. (2004); Nguyen and Tait (2023); Smith and Tait (2024); Servant and Tait (2003); Graham et al. (2016); Hebecker et al. (2024); Cyncynates and Weiner (2025a, b); Carenza et al. (2025a); Aprile and others (2019)
Discussion and conclusions — In this Letter, we investigate the ionization of neutral gas produced in dense MCs by decaying/annihilating DM. In particular, sub-keV DM could generate UV/X-ray photons that are able to produce a high and homogeneous ionization across the cloud since they are absorbed efficiently within such systems. In fact, we show that the ionization produced by the photon emission from sub-keV DM saturates the ionization that is observed in different MCs for decay lifetimes, or annihilation rates, that are not ruled out by any other constraint. Therefore, we test for the first time MCs as a target for indirect DM searches, demonstrating that they allow us to improve current constraints for decaying DM particles such as ALPs, dark photons, scalar particles or sterile neutrinos for masses from a few tens of eV to slightly below keV.
We begin by identifying local MCs as a particularly robust probe of this kind of DM. This is because estimating their ionization rate requires only knowledge of the local DM density without the need to assume a specific DM distribution model. Using the local MC L1551 we derive constraints that are similar, or even better, than the previous leading constraints. Then, we examine the case of G28.37+00.07, or the DRAGON cloud, which has been studied previously in detail and shows regions of it that are exceptionally quiescent in terms of their star formation activity, with ionization rates compatible with s-1. This constitutes what we call our fiducial constraint, and allows us improve previous constraints from CMB and LeoT gas heating up to several hundreds of eV. Finally, we use the case of G1.4-1.8+87, an MC pc away from the GC, to show the potential of using clouds near the GC. For this case, we take the values obtained in Ref. Bhoonah et al. (2019) to illustrate that, assuming an NFW DM density profile, we could reach the strongest DM bounds to date in a mass range from a few tens of eV to the keV. However, this must be taken as an optimistic forecast since the characterization of the temperature of that could was put in doubt in Ref. Farrar et al. (2020). Clouds that are closer to the GC can be found, therefore providing even stronger DM bounds. Finally, we highlight that the most promising targets for indirectly searching for sub-keV DM decays or annihilations into photons are IRDCs located near the GC but slightly above the Galactic plane. These regions are expected to be less affected by CR ionization, which is sourced in the Galactic disk. As a result, such clouds are likely to exhibit very low ionization levels, making any DM-induced effects more distinguishable. Although a more comprehensive survey of MCs will further refine the constraints presented here, our results already demonstrate the strong potential of this approach for probing different types of light DM particles.
However, even though these results are quite promising, one must keep in mind that the uncertainties in the measured ionization rates of Galactic molecular clouds are significant. These analyses require a combination of observations of molecular abundances and column densities to inform chemical models and infer the underlying ionization rates. Several factors contribute to these uncertainties. The chemistry of molecular clouds is highly sensitive to local conditions such as density, temperature, and radiation fields, which can vary across the cloud and are not always precisely known. Observational limitations, including instrumental sensitivity, line-of-sight integration, and assumptions about cloud geometry, further add to the uncertainty. Additionally, many studies report only best-fit values without fully quantified statistical errors, so the derived ionization rates often represent estimates rather than precise measurements. Despite these limitations, these analyses remain the most direct method to probe ionization processes in molecular clouds, and careful modeling allows researchers to account for uncertainties by considering conservative scenarios or quoting results as ranges. In this line, the James Webb Space Telescope has performed the first direct measurement of the ionization rate of Barnard 68 Bialy et al. (2025), which opens a window for refined ionization rate measurements in the next years.
We list a few expected future refinements below:
i) A more exhaustive survey of the measurements of dark MCs, dedicated to identify clouds with lower ionization fractions and closer to the GC. Such dedicated exploration and analysis could yield improvements of up to a few orders of magnitude in the derived DM constraints.
We show that dark clouds constitute a very interesting target for DM searches, hoping to increase the interest towards targeting dark MCs for observations.
iii) Probably the most obvious improvement would be that of modeling the ionization produced by CRs. Provided that CRs are expected to be the main ionization agent of dense MCs, including their contribution must certainly improve our constraints. Unfortunately, this estimation is not simple and requires an accurate knowledge of the distribution of CRs in different regions of the Galaxy and at very low energies, where our models are not robust enough to produce a reliable constraint. Reasonable estimates for local MCs already yield order-of-few improvements under conservative assumptions, as discussed above.
In conclusion, we have shown the potential impact of sub-keV DM decaying/annihilating into photons in dense gas regions, which can be specially relevant for probing ALPs and other DM particles. Our results demonstrate that astrochemical observations in MCs can serve as a novel and powerful probe for indirect searches of sub-keV DM. This approach offers particularly robust constraints across a range of particle physics models and has the potential to surpass existing bounds for DM masses between approximately eV and keV.
Acknowledgements — We thank Torsten Bringmann for fruitful discussion and Rebecca Leane for her useful comments and feedback. PDL has been supported by the Juan de la Cierva JDC2022-048916-I grant, funded by MCIU/AEI/10.13039/501100011033 European Union ”NextGenerationEU”/PRTR, and is currently supported by Ramón y Cajal RYC2024-048445-I grant, which is funded by MCIU/AEI/10.13039/501100011033 and FSE+. The work of PDL is also supported by the grants PID2021-125331NB-I00 and CEX2020-001007-S, both funded by MCIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”. PDL also acknowledges the MultiDark Network, ref. RED2022-134411-T. This project used computing resources from the National Academic Infrastructure for Supercomputing in Sweden (NAISS) under project NAISS 2024/5-666. PC and TTQN are supported by the Swedish Research Council under contract 2022-04283. This article/publication is based on the work from COST Action COSMIC WISPers CA21106, supported by COST (European Cooperation in Science and Technology).
Supplementary Material
Sub-keV dark matter can strongly ionize molecular clouds
Pedro De la Torre Luque Pierluca Carenza
Thong T. Q. Nguyen
SI Additional details
In this section, we provide additional figures that allow us to describe better our calculations. In the left panel of Fig. S1 we show the quantity J for several DM masses in a local MC like L1551. As expected, the lower the DM mass, the higher is J and, thus, the ionization rate. This behavior is due to the higher DM number density and the higher photoionization cross sections, which peaks around eV Yan et al. (2001). This can be also seen from the right panel of Fig. S1, where we show the H2 ionization rate caused by annihilating (dot-dashed line) and decaying (solid lines) DM in a local MC. The purple dashed line shows the level of ionization observed for L1551. We note that in this work we will constrain the annihilation rate into photons defined as , where is the total DM annihilation rate and is the branching ratio for photons in the final state. Similarly, we will always refer to DM decay into photons.
Another important factor in our study is the opacity of each cloud, which characterizes how efficiently photons can escape from the MC. The opacity is computed as . A very high opacity (i.e. an absorption probability close to unity) implies that photons cannot escape the molecular cloud and instead deposit all of their energy into the cloud in the form of ionization. As we see from Fig S2, while the high density of the P8 region in the DRAGON cloud makes it effectively opaque to photons with energies below 1 keV, in the L1151 and G1.4–1.8+87 clouds there exists an energy above which photons escape efficiently, producing little ionization above a few hundred keV.
Then, we show the uncertainties in our model-independent constraints on and for the DRAGON and G1.4-1.8+87 clouds associated to the assumed DM density profile. In particular, we take two extreme assumptions for the DM profile: the most pessimistic choice, which is the Burkert profile Burkert (1995) (Eq. 8) and the most optimistic choice, a Moore-like profile Moore and others (1999) (characterized by a contraction index ) (Eq. 7). Their parametric forms are the following:
For the Burkert profile (Eq. 8), we use a scale radius of kpc Cirelli et al. (2011), while for the Moore profile, we employ kpc. For the NFW profile, that is used as reference, kpc Di Mauro (2021). In all these cases the density normalization, is found by imposing that the local DM density is GeV/cm3. As we see in Fig. S3, the constraint obtained from the ionization in the DRAGON cloud is not very affected by the choice of DM density distribution, either for decay of annihilation, which makes this cloud an optimal target for robust DM searches. However, the projected constraint based on G1.4-1.8+87 varies significantly depending on the choice of DM distribution, making it less robust, but with tremendous potential under the assumption of a cuspy DM density distribution in the Galaxy.
SII Other particle models
@@
While the case of ALPs is probably the most interesting among current light DM candidates, in this appendix we explore the constraints for other particle models coupling to photons. First, we report the constraint for scalar particles as DM, that have mixing angle with the SM Higgs Dev et al. (2021); Nguyen et al. (2024). We calculate this mixing angle from the scalar dark matter decay to two photons lifetime as
@@
where is the fine structure constant, GeV is the electroweak vacuum expectation value and mS is the mass of the scalar particle. A comparison of our constraints with previous ones, for scalar DM, is shown in the left panel of Fig. S4. As for the case of ALPs, we are able to improve on existent constraints, such as those from CMB spectral distortions Bolliet et al. (2021a), LeoT gas heating Wadekar and Farrar (2021) or the model independent constraint on decaying DM (dDM) derived in Refs. Audren et al. (2014); Poulin et al. (2016), and our results are complementary with X-ray constraints Gewering-Peine et al. (2017); Essig et al. (2013); Foster et al. (2021), constraints from the CMB anisotropies Slatyer (2016) and stellar constraints Fiorillo et al. (2025); Bottaro et al. (2023). Interestingly, MCs can probe a region of the parameter space that cannot be currently tested from arguments of the stars dynamics.
@@
where m is the sterile neutrino mass and GF is the Fermi constant that represents the strength of the weak force. In this case, instead of producing two photons in the decay, only one is produced. Therefore, the value adopted is a factor of 2 lower than the two-photon final state. These bounds are shown in the right panel of Fig. S4, compared to previous ones. As one can see, the parameter space that is covered by MCs is already well tested by other mechanisms, specially the Lyman-alpha forest Bringmann et al. (2023). Also the Tremaine-Gunn constraint Tremaine and Gunn (1979) rules out sterile neutrinos with masses below keV. We report the optimistic constraint derived from G1.4-1.8+87, which may become relevant again if model-dependent approaches reopen this region of parameter space Dev et al. (2025).
@@
Finally, we have also calculated these constraints for the case of vector bosons decaying into three photons. This can be the case of dark photons or the B-L vector, which are popular candidates for DM particles. In this case, the spectrum of photons is given by a continuum spectrum, instead of a line-like spectrum. The photon flux at a given position can be expressed as as Pospelov et al. (2008); Linden et al. (2024a); McDermott et al. (2018); Cheung et al. (2025); Nguyen (2025); Redondo and Postma (2009); Krnjaic et al. (2023); Linden et al. (2024b)
@@
where . Using this spectrum, we compute Jχ→3γ for three-photon final state. The result is shown in Fig. S5, upper panel. A trend similar to the monochromatic injection appears, with a larger ionization at lower DM masses. In Fig. S5 (lower panel) we show the limit on the decay lifetime obtained for this injection of particles, for the case of L1551 and G1.4-1.8+87.
@@
We derive constraints on well-motivated vector dark matter models, such as the dark photon and the vector boson, using decay rate limits of the vector-to-three-photon final state. Both models arise from extensions of the Standard Model by anomaly-free gauge symmetries. The dark photon interacts with Standard Model particles via kinetic mixing with the photon field Holdom (1986); Fayet (1980, 1981); Kahn et al. (2017); Fabbrichesi et al. (2020); Caputo et al. (2021), while the vector boson couples directly to the Standard Model through a flavor-universal coupling Chun and Yun (2022); Carena et al. (2004); Nguyen and Tait (2023); Smith and Tait (2024). These symmetry extensions are extensively studied in the contexts of string theory, string compactification, and vector dark matter production via quantum fluctuations during inflation Servant and Tait (2003); Graham et al. (2016); Hebecker et al. (2024); Cyncynates and Weiner (2025a, b); Carenza et al. (2025a).
@@
Figure S6 presents the optimistic G1.4-1.8+87 constraints on vector dark matter models. We compute the decay width using the Euler-Heisenberg (EH) approximation, which is valid in the keV mass range McDermott et al. (2018); Linden et al. (2024b). The left panel shows the constraint on kinetic mixing , compared to limits from the universe’s age (4.6 Gyr), x-ray observations by INTEGRAL Linden et al. (2024b), and direct detection bounds from XENON Aprile and others (2019). For vector dark matter, we derive constraints on the flavor-universal coupling , assuming the decay proceeds dominantly via electron loops under the EH approximation. We compare this result with the INTEGRAL limit on the same three-photon final state Linden et al. (2024b) and with theoretical constraints based on neutrino decay channels and the universe’s lifetime Chun and Yun (2022).